water is not to spill, what is the minimum velocity the pail can have? (b) How much time per revolution is this equivalent to? 


1
Expert's answer
2021-10-31T18:12:59-0400

The minimum velocity "v" is defined by the conditions required for the water not to spill. This is equivalent to the following value of time per revolution:


"t=\frac 1n=\frac{2\pi}\omega=\frac{2\pi r}{v}."